How many ounces of pure water must be added to 80 ounces of a 35% salt solution to make a solution that is 14% salt
Answer by ankor@dixie-net.com(22740)

How many ounces of pure water must be added to 80 ounces of a 35% salt solution to make a solution that is 14% salt
:
let w = the amt of water required to do this
.14(w + 80) = .35(80)
.14w + 11.2 = 28
.14w = 28 - 11.2
.14w = 16.8
w = 16.8/.14
w = 120 oz of water required
:
:
Check this by finding the amt of salt in each mixture, should be unchanged
.35(80 = 28
.14(120+80) = 28